# Repetition Improves Language Model Embeddings

### v1

@@ Abstract
Recent approaches to improving the extraction of text embeddings from autoregressive large language models (LLMs) have largely focused on improvements to data, backbone pretrained language models, or improving task-differentiation via instructions. In this work, we address an architectural limitation of autoregressive models: token embeddings cannot contain information from tokens that appear later in the input. To address this limitation, we propose a simple approach, “echo embeddings,” in which we repeat the input twice in context and extract embeddings from the second occurrence. We show that echo embeddings of early tokens can encode information about later tokens, allowing us to maximally leverage high-quality LLMs for embeddings. On the MTEB leaderboard, echo embeddings improve over classical embeddings by over $9\%$ zero-shot and by around $0.7\%$ when fine-tuned. Echo embeddings with a Mistral-7B model achieve state-of-the-art compared to prior open source models that do not leverage synthetic fine-tuning data. [^1] Our code and pre-trained models are released at https://github.com/jakespringer/echo-embeddings.
Repetition Improves Language Model Embeddings
Jacob Mitchell Springer  Suhas Kotha Daniel Fried  Graham Neubig  Aditi Raghunathan Carnegie Mellon University {jspringe, suhask, dfried, gneubig, aditirag}@cs.cmu.edu
@@ 1 Introduction
In the recent past, the dominant pretrained language model paradigm for neural embeddings have been masked language models with bidirectional attention (Ni et al., 2021a; Raffel et al., 2020; Izacard et al., 2021; Wang et al., 2022; Jiang et al., 2022; Su et al., 2022; Xiao et al., 2023a; Li et al., 2023). However, more recent literature (Ma et al., 2023; Wang et al., 2023) has begun to scale these algorithms to modern autoregressive language models such as LLaMA-2 and Mistral (Touvron et al., 2023; Jiang et al., 2023a). Developing approaches to construct embeddings from autoregressive language models is promising: for many tasks, these models are the highest quality models available (Srivastava et al., 2022).
Figure 1: Conceptual overview of echo embeddings.
In this paper, we address a striking failure mode of autoregressive language models. This failure arises from the fact that for autoregressive language models, contextualized token embeddings—the vector of last-hidden-layer activations at the position of a particular input token—do not contain information from tokens that appear later in the sentence due to the causal attention mask. We demonstrate that such embeddings can fail to appropriately determine similarity when the early tokens are superficially similar but become dissimilar in important ways when using key information from end of the input.
We propose a strategy to overcome this limitation in autoregressive models through “echo embeddings.” With this approach, we repeat the inputs so that it appears twice in the context passed to the language model, and extract embeddings from the second occurrence. Repeating the input enables the contextualized token embeddings of the second occurrence of the passage to encode information from tokens that appear later in the passage by attending to their first occurrence in the passage. We show that echo embeddings do in fact allow embeddings of the early tokens to capture information about the later tokens.
We then evaluate echo embeddings on the standard Massive Text Embedding Benchmark (MTEB) leaderboard [^2] The MTEB leaderboard can be found at https://huggingface.co/spaces/mteb/leaderboard. In the zero-shot setting, echo embeddings improve on classical embeddings by over $9\%$ and provide consistent gains across all the different tasks for a variety of language models and scale. We then perform an apples-to-apples comparison when fine-tuning embeddings from Mistral-7B and continue to see consistent gains of echo over classical across the various tasks (by $0.7\%$ on average). Strikingly, echo embeddings with the strong Mistral-7B language model allows us to achieve state-of-the-art embedding quality, enabling autoregressive language models to match prior open-source top performing models that otherwise leveraged MLMs with bidirectional attention. [^3] The contemporaneous work by Wang et al. (2023) achieves state-of-the-art accuracy on MTEB using classical embeddings from autoregressive language models, but through finetuning on high quality synthetic data, which we believe is largely orthogonal to our contribution.
The approach of echo embeddings is conceptually well-motivated, extremely simple, and generally compatible with other innovations in extracting embeddings from autoregressive language models. As language models are likely to continue to improve over the coming years, echo embeddings can be a simple but powerful twist to classical embeddings that allow us to maximally leverage autoregressive language models.
@@ 2 Preliminaries
Our goal is to extract text embeddings that map a sentence $x$ to a vector $\phi(x)\in\mathbb{R}^{d}$ such that the semantic similarity between sentences is captured as similarity between their embeddings. In practice, we use the cosine similarity between embeddings to capture semantic similarity (detailed in Appendix B).
@@
As is standard, we extract embeddings from the activations of the final hidden layer. Each input token $x_{j}$ at position $j$ is associated with a contextualized token embedding which is the hidden layer representation $\phi_{j}(x)$.
We can pool the embeddings across all the tokens in different ways. In this work, we focus on two common strategies which have been considered by prior work (Reimers and Gurevych, 2019; Muennighoff, 2022; Zhang et al., 2023a; Wang et al., 2023).
A mean token embedding over a set of indices $A$, refers to the mean contextualized token embeddings at indices in $A:\phi_{A}(x)\coloneqq\frac{1}{\left|A\right|}\sum_{t\in A}\phi_{t}(x)$.
A last-token embedding refers to the contextualized token embedding of the last token in the input sequence, written $\phi_{-1}(x)$.
@@
Traditionally, embeddings are computed by simply passing the sentence to the model and extracting some pooling (e.g. mean or last-token) of the contextualized embeddings corresponding to the input sentence. We will refer to embeddings created in this way as “classical embeddings”. Additionally, one might first prompt the language model with an explanation of the task of interest followed by the sentence, and then pool the contextualized embeddings of the sentence tokens like before (Su et al., 2022).
3 Echo Embeddings
In this section, we first demonstrate a failure mode of classical embeddings, and motivate a new method that we call echo embeddings that addresses this failure.
3.1 Classical Embeddings Miss Bidirectional Information.
Sentence embeddings should aggregate information across the entire sentence. However, for autoregressive language models, the contextualized embedding at position $k$ $\phi_{k}(x)$ cannot encode information about tokens $x_{k+1},x_{k+2},\ldots$. Hence, the “meaning” encoded by the embeddings of tokens at the beginning of a sentence might inaccurately suggest they are similar (or dissimilar) to other tokens without considering the influence of tokens that come later. As a simple illustration, consider the following.
$q$ $\colon{{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[She loves summer]}}}{{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [but dislikes the heat]}}}$
$s^{-}$ $\colon{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[She loves summer]}}{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [for the warm evenings]}}$
$s^{+}$ $\colon{\color[rgb]{0.2,0.2,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.8}\text{[Summer is her favorite]}}{\color[rgb]{0.89,0.0,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.89,0.0,0.13}\text{ [but not the temp.]}}$
Here, the contextualized embeddings of the first half of $s^{+}$ and $s^{-}$ are both similar to $q$ because they do not attend to the second half of the sentence. As a result, the similarity between $q$ and $s^{-}$ would be overestimated by any pooling strategy that uses information from the first half. We address last-token pooling at the end of this section.
3.2 Echo Embeddings
We propose a simple fix to mitigate the failure above: we present the input sentence twice to the language model and extract contextualized embeddings from the second occurrence of the sentence. In principle, the contextualized embeddings of the second occurrence can attend to the entire sentence presented in the first occurrence. Furthermore, in order to encourage the second occurrence to actually “encode” information about the first, we instruct the language model to perform a generic task that requires using this information, e.g., “rewrite” or “repeat.”
Classical embeddings: Feed sentence $x$ to the language model and pool the contextualized embeddings of sentence $x$. Echo embeddings: Feed a prompt such as “Rewrite the sentence: $x$, rewritten sentence: xx” to the language model and pool the contextualized embeddings of the second occurence of $x$.
Key to our method is passing the sentence twice to the model and pool embeddings exclusively from the second occurrence. [^4] We find that minor variations of the echo embeddings prompt (e.g. change “rewrite” to “repeat”) work equally well and we provide an example list in Appendix B Other tricks from classical embeddings such as prompting the model with the downstream task of interest can be applied to echo embeddings as well.
@@ 3.2 Echo Embeddings
Figure 2: (left) We take the echo embeddings of only AA for query $q$ and sentences $s^{-},s^{+}$ and plot the distribution of cosine similarities, showing that echo embeddings encode later information in earlier tokens. (right) We plot the accuracy of classical and echo embeddings when the sentences have similar beginnings (Structure 1) and dissimilar beginnings (Structure 2).
3.3 Repetition Captures Bidirectional Info
In the previous section, we argued that classical embeddings suffer from the issue that contextualized embeddings of early tokens can miss out on information from the later tokens. But can simply repetition via echo embeddings solve this issue? We aim to test this by extracting embeddings from Mistral-7B on a simple controlled synthetic setting.
Given a query $q\colon[A,B]$ we construct sentence pairs $s^{+},s^{-}$ as follows. We make the first parts of each sentence identical to the query, but differ only in their second parts,
$q\colon[A,B]$; $s^{+}\colon[A,B^{+}]$; $s^{-}\colon[A,B^{-}]$,
where $B^{+}$ is $B$ but paraphrased and $B^{-}$ is semantically dissimilar to $B$. We query GPT-4 to generate examples of this structure. We describe the full procedure and prompts in Appendix B.

### v2

@@ Abstract
Bidirectional models are considered essential for strong text embeddings. Recent approaches to adapt autoregressive language models (LMs) into strong text embedding models have largely had the requirement to modify the LM architecture to be bidirectional. We challenge this premise by introducing “echo embeddings” which converts autoregressive LMs into high quality text embedding models without changing the architecture or requiring fine-tuning. By repeating the input and extracting embeddings from the repeated tokens—which have access to all original tokens—echo embeddings improve over classical LM embeddings by over 5% in zero-shot settings. Our zero-shot embeddings nearly match those obtained by bidirectionally-converted LMs that undergo additional masked-language modeling training. Echo embeddings are also compatible with supervised fine-tuning, matching or outperforming bidirectionally-converted LMs in an apples-to-apples comparison, even with an identical compute budget during training and inference. Overall, repetition is a simple and effective strategy to circumvent the need for bidirectional attention in embedding models, paving the way towards a unified architecture for all NLP tasks.
@@ 1 Introduction
Until recently, the dominant pretrained language model paradigm for neural embeddings have been masked language models with bidirectional attention (Ni et al., 2021a; Raffel et al., 2020; Izacard et al., 2021; Wang et al., 2022; Jiang et al., 2022; Su et al., 2022; Xiao et al., 2023a; Li et al., 2023). Very recent literature has made progress generating high quality embeddings from autoregressive language models. However, these approaches either require large-scale synthetic data (Wang et al., 2023) or for the architecture of the base model to be “cast” to have bidirectional attention (BehnamGhader et al., 2024; Muennighoff et al., 2024). The requirement to change the architecture prevents the extraction of high quality embeddings without additional training. Further, this requirement is an obstacle to the development of truly unified models that can generate text and produce embeddings simultaneously without changing the architecture between encoding text for embeddings and generating tokens.
Is bidirectional attention crucial for high quality embeddings? Intuitively, autoregressive models fail to capture information across the entire input due to the causal attention mask. We show this concretely in a toy setup in Section 3. In this work, we present a surprisingly effective way to overcome this limitation without requiring bidirectional attention or fine-tuning.
Our method, “echo embeddings” is simple: we prompt the model with the input twice and extract embeddings from the second occurrence of the input. Even with causal attention, the embeddings of the second occurrence can access the entire input in the first occurrence. The second key ingredient is to prompt the language model to repeat the input (or, for example, to fix, rephrase, or otherwise reconstruct the input). This simple change enables the model to extract bidirectional information. As a result, echo embeddings offer strong performance in the entirely zero-shot setting—without any additional fine-tuning.
In comparison with the classical method of extracting embeddings, the zero-shot performance of echo embeddings is around 5% higher. This nearly matches the performance of the recent more expensive method of casting to bidirectional attention followed with an addititional unsupervised MLM training (BehnamGhader et al., 2024). Even in comparison to the strong zero-shot baseline of PromptEOL (Jiang et al., 2023b)—a method in which the model is prompted to summarize the input in a single token—echo embeddings outperform by a significant margin, and is less sensitive to the exact choice of prompt. We emphasize that echo embeddings are simple, cheap, and flexible.
The current best embeddings still require additional fine-tuning with a supervised objective. We find that echo embeddings slightly outperform the classical approach to extracting embeddings, even when the architecture is cast to be bidirectional, and even when echo embeddings is given an identical compute budget. We speculate that these gains are because echo embeddings better preserve and leverage large-scale pretraining because there are no architectural changes. This is in contrast to the bidirectional relaxation approach, which requires the model to adjust its weights to perform well with bidirectional attention.
Echo embeddings come with a lot of benefits: simple to apply, works zero-shot without additional training, allows a unified model to be used for various tasks, and offers (slightly) better performance than casting to bidirectional attention offers. One apparent drawback is that repetition doubles the compute cost. To account for this, in all our experiments, we report “compute matched” results for both inference time and training time (when applicable). While the gains are now slightly smaller, we see that echo embeddings still perform well. In total, we believe that the echo embedding approach, by virtue of its simplicity and effectiveness, is an important step in bridging the gap between traditionally bidirectional embedding models and the current predominant paradigm of autoregressive language models.
@@ 2 Preliminaries
Our goal is to extract text embeddings that map a sentence $x$ to a vector $\phi(x)\in\mathbb{R}^{d}$ such that the semantic similarity between sentences is captured as similarity between their embeddings. In practice, we use the cosine similarity between embeddings as a similarity metric (detailed in Appendix C).
@@
As is standard, we extract embeddings from the activations of the final hidden layer. Each input token $x_{j}$ at position $j$ is associated with a contextualized token embedding which is the hidden layer representation $\phi_{j}(x)$. We can pool the embeddings across all the tokens in different ways. In this work, we focus on two common strategies which have been considered by prior work (Reimers & Gurevych, 2019; Muennighoff, 2022; Zhang et al., 2023a; Wang et al., 2023).
Pooling.
A mean token embedding over a set of indices $A$, refers to the mean contextualized token embeddings at indices in $A:\phi_{A}(x)\coloneqq\frac{1}{\left|A\right|}\sum_{t\in A}\phi_{t}(x)$. A last-token embedding refers to the contextualized token embedding of the last token in the input sequence, written $\phi_{-1}(x)$.
@@
Traditionally, embeddings are computed by simply passing the sentence to the model and extracting some pooling (e.g. mean or last-token) of the contextualized embeddings corresponding to the input sentence. We will refer to embeddings created in this way as “classical embeddings”. Additionally, one might first prompt the language model with an explanation of the task of interest followed by the sentence (Su et al., 2022).
3 Echo Embeddings from LLMs
@@
Figure 2: A. Accuracy of classical embeddings (mean and last-token pooling) and echo embeddings on the entire toy dataset. B. Accuracies of classical and echo embeddings split by the different structures of the data (S1 and S2, see text). C. We take the echo embeddings of only the first parts $A$ of query $q=[A,B]$ and sentences $s^{-},s^{+}$ and plot the distribution of cosine similarities, showing that echo embeddings encode can later information in earlier tokens.
In this section, we describe our general methodology to extract high quality embeddings from autoregressive language models without bidirectional attention. We start with a simple synthetic dataset that demonstrates why causal attention might inhibit embeddings from capturing information reliably across the entire context (Section 3.1). We then present our proposed method of Echo Embeddings in Section 3.2, and test that our simple and intuitive approach in fact addresses the failure mode identified in the synthetic setup without introducing bidirectional attention.
3.1 Failure of classical embeddings with causal attention
High-quality embeddings should effectively capture various text structures, including where similarity between inputs is strongly affected by either the early parts or the later parts of the inputs. We create a dataset with two “opposite” structures (S1 and S2) and show the classical embeddings with causal attention cannot capture both simultaneously.
The dataset we construct consists of a query $q$ and a similar sentence $s^{+}$ and a dissimilar sentence $s^{-}$. We measure the performance of embedding $\phi(\cdot)$ by computing the fraction where $\operatorname{Sim}(q,s^{+})>\operatorname{Sim}(q,s^{-})$. Each query and sentence is a concatenation of two strings, allowing us to vary whether early (or later) parts are similar (or discriminatory). Notationally, $A,A^{+}$, $B,B^{+}$ are semantically similar, while $A,A^{-}$, $B,B^{-}$ semantically dissimilar. We query GPT-4 to generate these various realistic but synthetic strings (see Section C).
S1 (Early discriminatory; late redundant). For a query $q=[A,B]$, in this structure, the early tokens provide information to discriminate between $s^{+},s^{-}$ while the later tokens are all similar. In other words, $s^{+}:[\mathbf{A^{+}},B^{+}]$ and $s^{-}:[\mathbf{A^{-}},B]$, where the discriminatory parts are bolded for convenience.
S2 (Early redundant; late discriminatory). Analogously, we consider a second “opposite” structure where the early tokens are all similar but the later tokens can discriminate between $s^{+},s^{-}$. For query $q=[A,B]$, we have $s^{+}:[A^{+},\mathbf{B^{+}}]$ and $s^{-}:[A,\mathbf{B^{-}}]$, where the discriminatory parts are bolded for convenience.
We show the results of zero-shot classical embeddings extracted from Mistral-7B-Instruct-v0.1 in Figure 2 (A) on a dataset that has a mixture of both structures. We see that neither mean nor last token pooling allows classical embeddings to achieve better-than-random performance.
To get some insight into the failure, we plot the performance on each substructure separately in Figure 2 (B). Last-token pooling disproportionately weights the later parts of the sentence thereby missing discriminatory information in the early tokens in S1. In general, last-token pooling does not work well in a zero-shot manner on real data as well, as we see in the next section.
Mean-token pooling is better suited to the zero-shot setting and can allow us to extract information from all parts of the input. This works well in structure S1. However, the causal attention produces a different issue: the contextual embedding at position $k$, $\phi_{k}(x)$ cannot encode information about the later tokens $x_{k+1},x_{k+2},\ldots$. This causes mean-token embeddings to exaggerate early token similarity and dilute discriminatory information in the later parts because of causal attention. This effect is especially exacerbated when the early tokens are similar, as in S2. In total, classical embeddings fail when the dataset contains inputs of both instances—where the discriminatory information could be in the early or later tokens. We show that this failure mode of classical embeddings can be overcome in a completely zero-shot manner using a simple method which we describe next.
3.2 Echo Embeddings: A General Framework
We saw in Section 3.1 that while mean-token pooling aggregates information across all tokens, the contextualized representations of early tokens are suboptimal because they cannot attend to the later tokens due to the causal attention. Replacing the causal with bidirectional attention is a natural solution, and common intuition seems to be that this is necessary. We propose a new general embedding framework—echo embeddings—that captures contextualized information from all tokens in the input with only causal attention:
Echo embeddings method: Prompt the language model to act as an autoencoder, e.g., by asking the model to {repeat, rephrase, fix, etc.} the input; feed the sentence $x$ to the language model twice; pool the contextualized embeddings of the second occurrence of $x$. Ex: “Rewrite the sentence: $x$; rewritten sentence: $\bf{x}$”; embeddings are pooled from the bolded $\bf{x}$.
The second occurrence can attend to everything, and hence extracting contextualized representations from the second occurrence could capture information from the entire input. Furthermore, in order to encourage the second occurrence to actually “encode” information about the first, we instruct the language model to perform a generic task that requires using this information, e.g., “rewrite” or “repeat”. The exact prompt, template, or task used to extract embeddings can be varied to suit the model and downstream task. We find empirically that the performance of echo embeddings are not particularly sensitive to the exact phrasing of the prompt (Section 4.2). We provide a list of example prompts in Appendix C.
While echo embeddings is simple and intuitive, does it actually effectively capture bidirectional information?
We apply echo embeddings with mean-token pooling on the synthetic dataset introduced in Section 3, and present the results in Figure 2. Echo embeddings with mean-token pooling works significantly better than classical embeddings and works well on both structures in the data—whether the discriminatory information is in the early or later tokens.